A) 0.4059 In each case show the result of calculations and circle one of the multiple choice answers. 7. A mountain search and rescue team receives an average of u-0.75 calls per day. Find B) 0.2706 C) 0.5412 D) 1.0824 the probability that on a randomly selected day, they will receive more then 1 call? Show the resultof calculations.and circle one of the multiple choice answers (6 points). 1534 B) 0.8266 C) 0.3542 D) 0.0.1734 乂 ,
Answer #1

Let, X be the number of calls receives by mountain search and rescue team per day.

X follows Poisson distribution with mean \mu = 0.75

We have to find P( X > 1 )

P( X > 1 ) = 1 - P( X <= 1 )

Using Excel function,   =POISSON( X , mean , 1 )

P( X <= 1 ) = POISSON( 1 , 0.75 , 1 ) = 0.826641

So, P( X > 1 ) = 1 - 0.826641 = 0.1734

Probability that on randomly selected day,they will receive more then 1 calls is 0.1734